Find the distance between the point and the line/plane 

a. (-1,4);x-3y+2=0

b. (-1,-1,2);2x+5y-6z=4


1
Expert's answer
2021-10-10T16:40:04-0400

Distance between point (x1, y1) and the line ax + by + c = 0 is (ax1 + by1 + c)/"\sqrt{(a^2+b^2)}"

Distance between point (–1, 4) and the line x – 3y + 2 = 0 is |(–1 – 3 × 4 + 2)/"\sqrt{(1^2+(-3)^2)}" |

                                                                                            = |–11/"\sqrt{10}" |

                                                                                            = 3.4785

 

 

Distance between point (x1, y1, z1) and the plane ax + by + cz + d = 0 is (ax1 + by1 + cz1)/"\sqrt{(a^2+b^2+c^2)}" )

Distance between point (–1, –1, 2) and the plane 2x + 5y –6z – 4 = 0 is |(2(–1) + 5(–1) – 6 × 2)/"\sqrt{(2^2+5^2+(-6)^2)}" |

                                                                                       = |–19/"\sqrt{65}" |

                                                                                       = 2.3567